2. A regular square pyramid has an altitude of 12 m and volume of 196 m3. Determine the slant height of the pyramid.
1
Expert's answer
2020-12-08T10:18:13-0500

Consider a regular square pyramid

Let "a=" the base edge, "h=" the altitude, and "s=" the slant height.

The volume "V" of the frustum of a regular square pyramid is


"V=\dfrac{1}{3}a^2h"

Then

"a=\sqrt{\dfrac{3V}{h}}=\sqrt{\dfrac{3(196)}{12}}=7 (m)"

By the Pythagorean Theorem from the right triangle


"s^2=(\dfrac{a}{2})^2+h^2"

"s=\sqrt{(\dfrac{a}{2})^2+h^2}=\sqrt{(\dfrac{7}{2})^2+(12)^2}=12.5(m)"

12.5 m